# Trigonometry of partially truncated hyperbolic triangles and tetrahedra

### v1

@@
Jason DeBlois
@@
The first main results of this note, proved in Section 2, establish additional such laws which are not already in [4] nor in Ratcliffe’s Foundations of hyperbolic manifolds [6]. Proposition 2.1 pertains to hyperbolic quadrilaterals, and 2.2 to pentagons, in which the non-truncated vertices are ideal (see Section 1 below). In these results, the length of a horospherical cross-section at an ideal vertex plays the role that would be played by the dihedral angle at an ordinary vertex.
In Section 3 we turn to the analogous situation in dimension three. Our second main result concerns truncated tetrahedra in $\mathbb{H}^{3}$, which are homeomorphic to affine $3$-simplices with open neighborhoods of their vertices removed. Proposition 3.5 bounds the transversal length of such a tetrahedron, meaning the minimum distance between a specified pair of opposite internal edges, in terms of the tetrahedron’s set of internal edge lengths. Here an edge is internal if it is not contained in a face resulting from truncation.
In both settings we follow the approach of [6, Ch. 3] and use the hyperboloid model for $\mathbb{H}^{n}$, where it is taken as a subset of $\mathbb{R}^{n+1}$ equipped with the Lorentzian inner product, a certain non positive-definite bilinear form. Vectors of the ambient $\mathbb{R}^{n+1}$ carry information about different objects of $\mathbb{H}^{n}$, depending on the sign of their self-pairing, and the Lorentzian inner product of two such vectors carries information about the hyperbolic distance between the objects the vectors encode. This property is exploited in [6] and also in previous work, eg. by Epstein–Penner [3]. We describe it in Section 1 and subsequently leverage it to encode partially truncated triangles using just three vectors, and truncated tetrahedra using just four, using the pairings of these vectors to prove our trigonometric formulas.
The results of this note are applied in producing lower bounds on volumes of hyperbolic $3$-manifolds with totally geodesic boundary, in other papers that have appeared or will soon.
@@ 1. Background: the meaning of vectors in the hyperboloid model
and $\mathbf{x}$ is said to be space-like, light-like, or time-like respectively as $\mathbf{x}\circ\mathbf{x}$ is positive, zero, or negative. The Lorentzian norm of $\mathbf{x}$ is $\|\mathbf{x}\|=\sqrt{\mathbf{x}\circ\mathbf{x}}$, where the square root is taken to be positive, zero, or positive imaginary in the respective cases above. The light cone is the set of light-like vectors, and its interior is the set of time-like vectors. A time-like or light-like vector is positive if its first entry is. The hyperboloid model $\mathbb{H}^{n}$ of hyperbolic space is the set of positive vectors with Lorentzian norm $i$ in $\mathbb{R}^{n+1}$, equipped with the distance $d_{H}$ defined by
@@
(We note that the following version of the Cauchy-Schwartz inequality follows from the usual one: for positive vectors $\mathbf{x}$ and $\mathbf{y}$ with $\mathbf{x}\circ\mathbf{x}\leq 0$ and $\mathbf{y}\circ\mathbf{y}\leq 0$, $\mathbf{x}\circ\mathbf{y}\leq-\sqrt{(\mathbf{x}\circ\mathbf{x})(\mathbf{y}\circ\mathbf{y})}$, with equality if and only if they are linearly dependent, see eg. formula (1.0.2) of [1].)
The distance function $d_{H}$ above is determined by the Riemannian metric on $\mathbb{H}^{n}$ given, at each $\mathbf{x}\in\mathbb{H}^{n}$, by restricting the Lorentzian inner product to $T_{\mathbf{x}}\mathbb{H}^{n}=\mathbf{x}^{\perp}\doteq\{\mathbf{v}\,|\,\mathbf{v}\circ\mathbf{x}=0\}$. (This restriction is positive-definite since $\mathbf{x}$ is time-like, see [6, Theorem 3.1.5].) The isometry group of $\mathbb{H}^{n}$ is the group $O^{+}(1,n)$ of matrices preserving the Lorentzian inner product and the sign of time-like vectors, see [6, §3.1], acting on $\mathbb{H}^{n}$ by restriction.
@@ 1.1. The meaning of light-like vectors
The horosphere determined by a positive light-like vector $\mathbf{x}\in\mathbb{R}^{n+1}$ is $S=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}=-1\}$. The horoball bounded by $S$ is the set $B=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}\geq-1\}$.
@@
To see this, set $F(\mathbf{v})$ equal to $\mathbf{u}_{1}$ and solve for $\mathbf{v}\circ\mathbf{v}$ by taking the Lorentzian inner product of both sides with $\mathbf{u}_{0}$. Using the formula for $d_{H}(\mathbf{u}_{0},\mathbf{u}_{1})$ given above we obtain the comparison equation $d_{S}(\mathbf{u}_{0},\mathbf{u}_{1})/2=\sinh(d_{H}(\mathbf{u}_{0},\mathbf{u}_{1})/2)$. We note that this implies in particular that the isometric embedding $F$ is proper; that is, $S$ has compact intersection with any compact set of $\mathbb{H}^{n}$.
@@
For the parametrized curve $\gamma$ defined in the statement, direct computation reveals that $\gamma(t)\circ\gamma(t)=-1$ for all $t$, so $\gamma$ maps into $\mathbb{H}^{n}$, and that $\gamma^{\prime\prime}(t)=\gamma(t)$. Therefore $\gamma$ is a hyperbolic geodesic, by [6, Theorem 3.2.4]. More direct computation shows that $\gamma(0)=\mathbf{v}$ and $\gamma(d)$ is the nearest point $\mathbf{u}$ to $\mathbf{v}$ on $S$ described in (2). ∎
@@
Note that $\gamma(0)=\mathbf{u}\in S_{1}$ and $\gamma(d)\in S_{0}$. The more-symmetric formula given in the statement is obtained by translating the parametrization, replacing $t$ by $t-d/2$.
@@ 1.2. The meaning of space-like vectors
As defined in [6, §3.2], a hyperplane of $\mathbb{H}^{n}$ is its intersection with a time-like, codimension-one vector subspace of $\mathbb{R}^{n+1}$. Corollary 4 of [6, §3.2] implies that the group of hyperbolic isometries acts transitively on the set of hyperplanes. Thus each hyperplane is the polar hyperplane to a space-like vector, since for instance $(\mathbb{R}^{n}\times\{0\})\cap\mathbb{H}^{n}$ is the polar hyperplane to $\mathbf{e}_{n+1}=(0,\ldots,0,1)$.
@@
where the sign of “$\pm$” above is negative if $\mathbf{v}_{1}$ belongs to the half-space $H_{2}$ bounded by $P_{2}$ with $\mathbf{y}_{2}$ as outward unit normal vector, and negative otherwise.
@@ 2. Dimension two
[…]
@@ 3. Dimension three: transversals of truncated tetrahedra
Recall the standard fact, proved in eg. [2, Lemma 2.3], that for any collection of three disjoint planes in $\mathbb{H}^{3}$ there is a unique fourth plane meeting each of the original three at right angles. The next result identifies a key half-space bounded by such a plane, in the present setting.
@@
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$. For any fixed $i$, let $\widehat{P}_{i}$ be the plane that intersects $P_{j}$ at right angles for each $j\neq i$. If $P_{i}$ does not meet $\widehat{P}_{i}$ orthogonally then there is a single half-space $\widehat{H}_{i}$ bounded by $\widehat{P}_{i}$ such that for all $j\neq i$, $\widehat{H}_{i}$ contains the shortest geodesic arc from $P_{j}$ to $P_{i}$.
@@
Fix $i\in\{1,2,3,4\}$, and let us establish some notation. For each $j\in\{1,2,3,4\}$ let $\mathbf{y}_{j}$ be an outward unit normal, in the sense described below Lemma 1.5, to the half-space bounded by $P_{j}$ that contains each other $P_{j^{\prime}}$. For any $j\neq j^{\prime}$, $\mathbf{y}_{j}$ and $\mathbf{y}_{j^{\prime}}$ are then oppositely-oriented tangent vectors to the hyperbolic geodesic intersecting $P_{j}$ and $P_{j^{\prime}}$ perpendicularly, so $\mathbf{y}_{j}\circ\mathbf{y}_{j^{\prime}}<0$ by Lemma 1.8. For each $j\neq i$, let $\mathbf{v}_{j}$ be the point of intersection between $P_{j}$ and the geodesic intersecting it and $P_{i}$ perpendicularly. For each $j$, Lemma 1.9 gives:
@@
We now consider the complementary case to that of Lemma 3.1, for planes $P_{1},P_{2},P_{3},P_{4}$ satisfying its hypotheses: if there exists an $i$ such that $P_{i}$ intersects $\widehat{P}_{i}$, defined as in the Lemma, at right angles, then the single plane $\widehat{P}\doteq\widehat{P}_{i}$ intersects all four planes at right angles and thus also equals $\widehat{P}_{j}$ for each $j\neq i$.
@@
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$. For each $i<j\in\{1,2,3,4\}$, denote the shortest arc in $\mathbb{H}^{3}$ joining $P_{i}$ to $P_{j}$ as $\lambda_{ij}$.
If $\widehat{P}_{i}$ as in Lemma 3.1 does not meet $P_{i}$ orthogonally for any $i$, the truncated tetrahedron determined by the $P_{i}$ is
@@
If $\widehat{P}_{i}$ does meet $P_{i}$ orthogonally for some $i$, then taking $\widehat{P}=\widehat{P}_{i}$ to be the unique plane that intersects each $P_{i}$ at right angles and renumbering the $P_{i}$ so that the geodesic containing $\lambda_{13}$ separates $P_{2}\cap\widehat{P}$ from $P_{4}\cap\widehat{P}$, we define $\Delta$ as a degenerate truncated tetrahedron by:
@@
where $h_{12}$ is the half-plane in $\widehat{P}$ that is bounded by the geodesic containing $\lambda_{12}$ and contains $P_{3}\cap\widehat{P}$ and $P_{4}\cap\widehat{P}$; and so on.
For each $i<j\in\{1,2,3,4\}$, call $\lambda_{ij}$ an internal edge of $\Delta$. The internal edge opposite $\lambda_{ij}$ is $\lambda_{kl}$, where $k<l\in\{1,2,3,4\}-\{i,j\}$. For each $i$, the internal face opposite $P_{i}$ is the right-angled hexagon in $\Delta\cap\widehat{P}_{i}$ bounded by the internal edges $\lambda_{jk}$, for each pair $j<k\in\{1,2,3,4\}-\{i\}$, and arcs of the $P_{j}$, $j\neq i$. The non-internal faces and edges of $\Delta$ are external. Each of these is entirely contained in $P_{i}$ for some $i$.
@@
[…]
@@
The main results of this section record some observations about the lengths of transversals of truncated tetrahedra. Before embarking on this we record the following basic geometric fact.
@@
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$; and for any $i<j\in\{1,2,3,4\}$, let $\tilde{\lambda}_{ij}$ be the geodesic intersecting $P_{i}$ and $P_{j}$ at right angles. Fixing such an $i<j$ and $k<l\in\{1,2,3,4\}-\{i,j\}$, and fixing parametrizations $\tilde{\lambda}_{ij}(s)$ and $\tilde{\lambda}_{kl}(t)$ by arclength, the function $D(s,t)$ that records the hyperbolic cosine of the distance from $\tilde{\lambda}_{ij}(s)$ to $\tilde{\lambda}_{kl}(t)$ has a unique critical point in $\mathbb{R}^{2}$, at which it attains an absolute minimum. The absolute minimum value depends only on the pairwise distances between the planes.
All assertions above but the final one follow from standard geometric facts in a straightforward way, and the last is implied by [5, Prop. 2.7]. Here we give a unified proof, following the general perspective taken in this note, which will set us up to prove this section’s main result.

### v2

@@
Jason DeBlois Department of Mathematics, University of Pittsburgh jdeblois@pitt.edu
@@
The first main results of this note, proved in Section 2, establish additional such laws which are not already in [7] nor other standard references that we know (eg. by Beardon [1] or Ratcliffe [10]). Proposition 2.1 pertains to hyperbolic quadrilaterals, and 2.2 to pentagons, in which the non-truncated vertices are ideal (see Section 1 below). In these results, the length of a horospherical cross-section at an ideal vertex plays the role that would be played by the dihedral angle at an ordinary vertex, and the distance to the cross-section plays the role of edge length.
In Sections 3 and 4 we turn to the analogous situation in dimension three. The results of Section 3 concern truncated tetrahedra in $\mathbb{H}^{3}$, which are homeomorphic to affine $3$-simplices with open neighborhoods of their vertices removed. Proposition 3.5 bounds the transversal length of such a tetrahedron, meaning the minimum distance between a specified pair of opposite internal edges, in terms of the tetrahedron’s set of internal edge lengths. Here an edge is internal if it is not contained in a face resulting from truncation. Section 4 addresses transversal lengths of partially truncated tetrahedra whose non-truncated vertices are ideal, successively treating the cases of one, two, three, and finally, in Proposition 4.8, four ideal vertices.
In all settings we follow the approach of [10, Ch. 3] and use the hyperboloid model for $\mathbb{H}^{n}$, where it is taken as a subset of $\mathbb{R}^{n+1}$ equipped with the Lorentzian inner product, a certain non positive-definite bilinear form. Vectors of the ambient $\mathbb{R}^{n+1}$ carry information about different objects of $\mathbb{H}^{n}$, depending on the sign of their self-pairing, and the Lorentzian inner product of two such vectors carries information about the hyperbolic distance between the objects the vectors encode. This property is exploited in [10] and also in previous work, eg. by Epstein–Penner [6]. We describe it in Section 1 and subsequently leverage it to encode partially truncated triangles using just three vectors, and truncated tetrahedra using just four, using the pairings of these vectors to prove our trigonometric formulas.
The results of Section 3 here are applied in [5] as one tool contributing to that paper’s lower bounds on volumes of hyperbolic $3$-manifolds with totally geodesic boundary. Similarly, we expect the results of Sections 2 and 4 to be used in forthcoming work on non-compact such manifolds. We are not aware of other references in the literature for the main results of Sections 2, 3, and 4. Section 2 of [8] also considers the geometry of partially truncated tetrahedra, but tracks them using dihedral angles and is focused on different questions; eg. existence and moduli.
@@ 1. Background: the meaning of vectors in the hyperboloid model
and $\mathbf{x}\neq\mathbf{0}$ is said to be space-like, light-like, or time-like respectively as $\mathbf{x}\circ\mathbf{x}$ is positive, zero, or negative. The Lorentzian norm of $\mathbf{x}$ is $\|\mathbf{x}\|=\sqrt{\mathbf{x}\circ\mathbf{x}}$, where the square root is taken to be positive, zero, or positive imaginary in the respective cases above. The light cone is the set of light-like vectors, and its interior is the set of time-like vectors. A time-like or light-like vector is positive if its first entry is. We note that the following version of the Cauchy-Schwartz inequality follows from the usual one, see eg. formula (1.0.2) of [3]:
Fact 1.1.
For positive vectors $\mathbf{x}$ and $\mathbf{y}$ with $\mathbf{x}\circ\mathbf{x}\leq 0$ and $\mathbf{y}\circ\mathbf{y}\leq 0$, $\mathbf{x}\circ\mathbf{y}\leq-\sqrt{(\mathbf{x}\circ\mathbf{x})(\mathbf{y}\circ\mathbf{y})}$, with equality if and only if they are linearly dependent.
The hyperboloid model $\mathbb{H}^{n}$ of hyperbolic space is the set of positive vectors with Lorentzian norm $i$ in $\mathbb{R}^{n+1}$, equipped with the distance $d_{H}$ defined by
@@
It follows from Fact 1.1 that this formula is well-defined. It is the distance function determined by the Riemannian metric on $\mathbb{H}^{n}$ given, at each $\mathbf{x}\in\mathbb{H}^{n}$, by restricting the Lorentzian inner product to $T_{\mathbf{x}}\mathbb{H}^{n}=\mathbf{x}^{\perp}\doteq\{\mathbf{v}\,|\,\mathbf{v}\circ\mathbf{x}=0\}$. (This restriction is positive-definite since $\mathbf{x}$ is time-like, see [10, Theorem 3.1.5].) The isometry group of $\mathbb{H}^{n}$ is the group $O^{+}(1,n)$ of matrices preserving the Lorentzian inner product and the sign of time-like vectors, see [10, §3.1], acting on $\mathbb{H}^{n}$ by restriction.
@@ 1.1. The meaning of light-like vectors
The horosphere determined by a positive light-like vector $\mathbf{x}\in\mathbb{R}^{n+1}$ is $S=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}=-1\}$. The horoball bounded by $S$ is the set $B=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}\geq-1\}$. We say that the projective class $[\mathbf{x}]$ of $\mathbf{x}$ is the ideal point of $S$ or of $B$.
@@
To see this, set $F(\mathbf{v})$ equal to $\mathbf{u}_{1}$ and solve for $\mathbf{v}\circ\mathbf{v}$ by taking the Lorentzian inner product of both sides with $\mathbf{u}_{0}$. Using the formula for $d_{H}(\mathbf{u}_{0},\mathbf{u}_{1})$ given above we obtain the comparison equation $d_{S}(\mathbf{u}_{0},\mathbf{u}_{1})/2=\sinh(d_{H}(\mathbf{u}_{0},\mathbf{u}_{1})/2)$. This implies in particular that the isometric embedding $F$ is proper; that is, $S$ has compact intersection with any compact set of $\mathbb{H}^{n}$.
@@
Remark.
For any $t\in\mathbb{R}$, $\gamma^{\prime}(t)=-\gamma(t)+e^{-d}e^{t}\mathbf{x}$ is a linear combination of $\gamma(t)$ and $\mathbf{x}$. Therefore by the discussion below Definition 1.2 it is normal to the horosphere through $\gamma(t)$ with ideal point $[\mathbf{x}]$.
@@
For the parametrized curve $\gamma$ defined in the statement, direct computation reveals that $\gamma(t)\circ\gamma(t)=-1$ for all $t$, so $\gamma$ maps into $\mathbb{H}^{n}$, and that $\gamma^{\prime\prime}(t)=\gamma(t)$. Therefore $\gamma$ is a hyperbolic geodesic, by [10, Theorem 3.2.4]. More direct computation shows that $\gamma(0)=\mathbf{v}$ and $\gamma(d)$ is the nearest point $\mathbf{u}$ to $\mathbf{v}$ on $S$ described in (2). ∎
@@
Note that $\gamma(0)=\mathbf{u}\in S_{1}$ and $\gamma(d)\in S_{0}$. The more-symmetric formula given in the statement is obtained by translating the parametrization, replacing $t$ by $t+d/2$.
@@ 1.2. The meaning of space-like vectors
As defined in [10, §3.2], a hyperplane of $\mathbb{H}^{n}$ is its intersection with a time-like, codimension-one vector subspace of $\mathbb{R}^{n+1}$. Corollary 4 of [10, §3.2] implies that the group of hyperbolic isometries acts transitively on the set of hyperplanes. Thus each hyperplane is the polar hyperplane to a space-like vector, since for instance $(\mathbb{R}^{n}\times\{0\})\cap\mathbb{H}^{n}$ is the polar hyperplane to $\mathbf{e}_{n+1}=(0,\ldots,0,1)$.
@@
Remark.
In the complementary case to Lemma 1.8 in which $\mathbf{x}$ as above lies in the ideal boundary of $P$, then for any $\mathbf{v}\in P$ the entire geodesic $\gamma$ from $\mathbf{v}$ in the direction of $\mathbf{x}$ from Lemma 1.3 lies in $P$. Thus $P$ contains points at arbitrarily small signed distance from $B$; in particular, it intersects it.
@@
where the sign of “$\pm$” above is negative if $\mathbf{v}_{1}$ belongs to the half-space $H_{2}$ bounded by $P_{2}$ with $\mathbf{y}_{2}$ as outward normal vector, and negative otherwise.
@@ 2. Dimension two
$\cosh\ell_{i}=\frac{e^{a_{i}}\cosh d+e^{a_{1-i}}}{e^{a_{i}}\sinh d}\quad\mbox{for}\ i=0,1.$
@@ 3. Dimension three: transversals of truncated tetrahedra
It is a standard fact, proved in eg. [4, Lemma 2.3], that for any collection of three disjoint planes in $\mathbb{H}^{3}$ there is a unique fourth plane meeting each of the original three at right angles. The next result re-establishes this in the present setting, for completeness, and it identifies a key half-space bounded by such a plane.
@@
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$. For any fixed $i$, there is a unique plane $\widehat{P}_{i}$ that intersects $P_{j}$ at right angles for each $j\neq i$. If $P_{i}$ does not meet $\widehat{P}_{i}$ orthogonally then there is a single half-space $\widehat{H}_{i}$ bounded by $\widehat{P}_{i}$ such that for all $j\neq i$, $\widehat{H}_{i}$ contains the shortest geodesic arc from $P_{j}$ to $P_{i}$.
@@
For each $j\in\{1,2,3,4\}$ let $\mathbf{y}_{j}$ be an outward unit normal, in the sense described below Lemma 1.6, to the half-space bounded by $P_{j}$ that contains each other $P_{j^{\prime}}$. As observed in the proof of Lemma 1.10, for any $j\neq j^{\prime}$, $\mathbf{y}_{j}$ and $\mathbf{y}_{j^{\prime}}$ are tangent vectors to the hyperbolic geodesic intersecting $P_{j}$ and $P_{j^{\prime}}$ perpendicularly, which has the form $\mathrm{Span}\{\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}\}\cap\mathbb{H}^{n}$. By their choices they are oppositely-oriented, so $\mathbf{y}_{j}\circ\mathbf{y}_{j^{\prime}}<0$ by Lemma 1.9.
For a fixed $i\in\{1,2,3,4\}$, let $\widehat{P}_{i}$ be the intersection with $\mathbb{H}^{3}$ of the span of the set of $\mathbf{y}_{j}$ for $j\neq i$. By the above, $\widehat{P}_{i}$ contains the mutual perpendicular geodesic to $P_{j}$ and $P_{j^{\prime}}$, for any such distinct $j$ and $j^{\prime}$, so it meets each such $P_{j}$ at right angles. Conversely, any plane $\widehat{P}$ that intersects any such $P_{j}$ at right angles contains its normal $\mathbf{y}_{j}$, since this is the tangent vector to a geodesic in $\widehat{P}$ that is perpendicular to $P_{j}$ at its point of intersection with $P_{j}$. Therefore $\widehat{P}=\widehat{P}_{i}$.
For each $j\neq i$, let $\mathbf{v}_{j}$ be the point of intersection between $P_{j}$ and the geodesic intersecting it and $P_{i}$ perpendicularly. For each $j$, Lemma 1.10 gives:
@@
We now consider the complementary case to that of Lemma 3.1, for planes $P_{1},P_{2},P_{3},P_{4}$ satisfying its hypotheses: if there exists an $i$ such that $P_{i}$ intersects $\widehat{P}_{i}$, defined as in the Lemma, at right angles, then the single plane $\widehat{P}\doteq\widehat{P}_{i}$ intersects all four planes at right angles and thus also equals $\widehat{P}_{j}$ for each $j\neq i$. The definition below accommodates both cases.
@@
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$. If for some (hence all) $i$, $\widehat{P}_{i}$ as in Lemma 3.1 does not meet $P_{i}$ orthogonally for any $i$, the truncated tetrahedron determined by the $P_{i}$ is
@@
If $\widehat{P}_{i}$ does meet $P_{i}$ orthogonally for some $i$, then taking $\widehat{P}=\widehat{P}_{i}$ to be the unique plane that intersects each $P_{i}$ at right angles, and renumbering the $P_{i}$ so that the perpendicular geodesic to $P_{1}$ and $P_{3}$ separates $P_{2}\cap\widehat{P}$ from $P_{4}\cap\widehat{P}$, we define $\Delta$ as a degenerate truncated tetrahedron by:
@@
where $h_{12}$ is the half-plane in $\widehat{P}$ that is bounded by the perpendicular geodesic to $P_{1}$ and $P_{2}$ that contains $P_{4}\cap\widehat{P}$ (hence also $P_{3}\cap\widehat{P}$); and so on.
For each $i<j\leq 4$, denote the shortest arc in $\mathbb{H}^{3}$ joining $P_{i}$ to $P_{j}$ as $\lambda_{ij}$ and call it an internal edge of $\Delta$. The internal edge opposite $\lambda_{ij}$ is $\lambda_{kl}$, where $k<l\in\{1,2,3,4\}-\{i,j\}$. For each $i$, the internal face opposite $P_{i}$ is the right-angled hexagon in $\Delta\cap\widehat{P}_{i}$ bounded by the internal edges $\lambda_{jk}$, for each pair $j<k\in\{1,2,3,4\}-\{i\}$, and arcs of the $P_{j}$, $j\neq i$. The non-internal faces and edges of $\Delta$ are external. Each of these is entirely contained in $P_{i}$ for some $i$.
@@
$\mathbf{v}_{1}$$\mathbf{y}_{3}$$\mathbf{y}_{2}$$\mathbf{y}_{4}$$\mathbf{y}_{1}$$\mathbf{v}_{3}$
@@
The main results of this section record some observations about the lengths of transversals of truncated tetrahedra. Before embarking on this we record the following basic calculus fact.
@@
Any positive linear combination $D(s,t)=C_{++}e^{s}e^{t}+C_{+-}e^{s}e^{-t}+C_{-+}e^{-s}e^{t}+C_{--}e^{-s}e^{-t}$ of the functions $e^{s}e^{t}$, $e^{s}e^{-t}$, $e^{-s}e^{t}$, and $e^{-s}e^{-t}$ is strictly convex and has a unique critical point $(s_{0},t_{0})$, at which it attains an absolute minimum. It satisfies:
(6) $e^{s_{0}}=\left(\frac{C_{-+}C_{--}}{C_{+-}C_{++}}\right)^{1/4},$ $e^{t_{0}}=\left(\frac{C_{+-}C_{--}}{C_{-+}C_{++}}\right)^{1/4},$ $\frac{D(s_{0},t_{0})}{2}=\sqrt{C_{++}C_{--}}+\sqrt{C_{+-}C_{-+}}.$
Furthermore, if $0<C_{++}<C_{+-},C_{-+}<C_{--}$ then $(s_{0},t_{0})\in(0,\infty)^{2}$.
@@
The Hessian of each of $e^{s}e^{t}$ and $e^{-s}e^{-t}$ is a scalar multiple of $\left(\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}\right)$, and that of each of $e^{s}e^{-t}$ and $e^{-s}e^{t}$ is a multiple of $\left(\begin{smallmatrix}1&-1\\ -1&1\end{smallmatrix}\right)$. Each of these matrices has 00 and $2$ as eigenvalues and $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ and $\left(\begin{smallmatrix}1\\ -1\end{smallmatrix}\right)$ as eigenvectors, but the null eigenvector of either is associated to the eigenvalue $2$ of the other. It follows that $D(s,t)$, being a positive linear combination of the four functions, has positive-definite Hessian and therefore is strictly convex.
We note that $D(s,t)\to\infty$ as $(s,t)\to\infty$. One can see this using the first summand $C_{++}e^{s}e^{t}$ for $(s,t)$ in the first quadrant; the summand $C_{-+}e^{-s}e^{t}$ for $(s,t)$ in the second; $C_{--}e^{-s}e^{-t}$ for $(s,t)$ in the third; and $C_{+-}e^{s}e^{-t}$ for $(s,t)$ in the fourth. $D(s,t)$ therefore attains an absolute minimum at a critical point, which is unique by convexity. To identify the critical point of $D(s,t)$ as defined above, we compute:
$\partial_{s}D(s,t)$ $=C_{++}e^{s}e^{t}+C_{+-}e^{s}e^{-t}-C_{-+}e^{-s}e^{t}-C_{--}e^{-s}e^{-t}$
$\partial_{t}D(s,t)$ $=C_{++}e^{s}e^{t}-C_{+-}e^{s}e^{-t}+C_{-+}e^{-s}e^{t}-C_{--}e^{-s}e^{-t}.$
Setting each of these equal to 00, subtracting bottom from top, and multiplying by $e^{s_{0}}e^{t_{0}}/2$ yields $0=C_{+-}e^{2s_{0}}-C_{-+}e^{2t_{0}}$, giving $e^{2s_{0}}=e^{2t_{0}}C_{-+}/C_{+-}$ for the critical point $(s_{0},t_{0})$. Adding the two equations and multiplying through by $e^{s_{0}}e^{t_{0}}/2$ yields $0=C_{++}e^{2s_{0}}e^{2t_{0}}-C_{--}$. After substituting for $e^{2s_{0}}$ we solve for $e^{2t_{0}}$, then plug the result back in to obtain the formulas of (6).
The sequence of inequalities in this result’s statement that starts with $0<C_{++}$ implies that $C_{--}/C_{++}$ is greater than both $C_{+-}/C_{-+}$ and its reciprocal $C_{-+}/C_{+-}$, which in turn implies that the quantities in parenthesis in (6) are each greater than $1$. Therefore if these inequalities hold then each of $s_{0}$ and $t_{0}$ is greater than 00. ∎
Lemma 3.4.
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$, and let $\Delta$ be the truncated tetrahedron determined by the $P_{i}$ as in Definition 3.2. For any fixed $i<j\in\{1,2,3,4\}$ and $k<l\in\{1,2,3,4\}-\{i,j\}$, let $\tilde{\lambda}_{ij}$ and $\tilde{\lambda}_{kl}$ be the geodesics respectively intersecting $P_{i}$ to $P_{j}$, and $P_{k}$ and $P_{l}$, at right angles. Fixing parametrizations $\tilde{\lambda}_{ij}(s)$ and $\tilde{\lambda}_{kl}(t)$ by arclength:
  1. (1)
    the function $D(s,t)$ that records the hyperbolic cosine of the distance from $\tilde{\lambda}_{ij}(s)$ to $\tilde{\lambda}_{kl}(t)$ has a unique critical point $(s_{0},t_{0})$ in $\mathbb{R}^{2}$, at which it attains an absolute minimum;
  2. (2)
    the absolute minimum value $D(s_{0},t_{0})$ depends only on the pairwise distances $\ell_{i^{\prime}j^{\prime}}$ between the $P_{i^{\prime}j^{\prime}}$, for $i^{\prime}\neq j^{\prime}\in\{1,2,3,4\}$;
  3. (3)
    the points $\tilde{\lambda}_{ij}(s_{0})$ and $\tilde{\lambda}_{kl}(t_{0})$ lie in $\Delta$, so $D(s_{0},t_{0})$ is the hyperbolic cosine of a transversal length of $\Delta$; and
  4. (4)
    written as $T(x,y;a,b,c,d)$, where $x=\cosh\ell(\lambda_{ij})$, $y=\cosh\ell(\lambda_{kl})$, and $a$, $b$, $c$, $d$ are hyperbolic cosines of the other four internal edge lengths of $\Delta$, this transversal length is invariant under even involutions of $\{a,b,c,d\}$, and strictly increasing in each of these variables, for any fixed $x,y>1$.
Convexity of the distance between two parametrized geodesics is well-known in broad generality (see eg. [2, Prop. 2.2] for the CAT(0) context). Conclusion (3) above is implied by [8, Prop. 2.7]. However our unified proof of all conclusions, which follows the general perspective taken in this note, will set us up to prove this section’s main result.
Proof.